### v1

# AdaWorld: Learning Adaptable World Models with Latent Actions

Abstract
World models aim to learn action-controlled prediction models and have proven essential for the development of intelligent agents. However, most existing world models rely heavily on substantial action-labeled data and costly training, making it challenging to adapt to novel environments with heterogeneous actions through limited interactions. This limitation can hinder their applicability across broader domains. To overcome this challenge, we propose AdaWorld, an innovative world model learning approach that enables efficient adaptation. The key idea is to incorporate action information during the pretraining of world models. This is achieved by extracting latent actions from videos in a self-supervised manner, capturing the most critical transitions between frames. We then develop an autoregressive world model that conditions on these latent actions. This learning paradigm enables highly adaptable world models, facilitating efficient transfer and learning of new actions even with limited interactions and finetuning. Our comprehensive experiments across multiple environments demonstrate that AdaWorld achieves superior performance in both simulation quality and visual planning.
Interactive Simulation, World Model, Latent Action
1 Introduction
Figure 1: Different world model learning paradigms. Previous methods require substantial action labels and expensive training to achieve action controllability in new environments. To overcome this limitation, we introduce latent actions as a unified condition for action-aware pretraining from videos, enabling highly adaptable world modeling. Our world model, dubbed AdaWorld, can readily transfer actions across contexts without training. By initializing action embeddings with the corresponding latent actions, AdaWorld can also be efficiently adapted into specialized world models through limited interactions and finetuning steps.
Intelligent agents should perform effectively across various tasks (Reed et al., 2022; Lee et al., 2022; Durante et al., 2024; Raad et al., 2024). A promising solution to this objective is developing world models that can simulate different environments (Wu et al., 2023, 2024; Yang et al., 2024c; Hansen et al., 2024). Recent world models are typically initialized from pretrained video models (Xiang et al., 2024; Agarwal et al., 2025; He et al., 2025). Despite improved generalization, these models still require substantial action labels and high training costs to acquire precise action controllability. While it is possible to annotate videos with pseudo labels for pretraining (Baker et al., 2022; Zhang et al., 2022), defining a unified action format for general environments is challenging. As a result, existing methods often require costly training when adapting to new environments with varying action specifications (Gao et al., 2024; Chi et al., 2024; Che et al., 2025). These limitations pose great challenges for transferring and learning new actions based on limited interactions and finetuning.
As humans, we can estimate the effects of different actions through limited experiences (Ha & Schmidhuber, 2018). This ability likely arises from our internal representations of actions learned from extensive observations (Rizzolatti et al., 1996; Romo et al., 2004; Dominici et al., 2011). These common knowledge can be reused across different contexts and associated with specific action spaces efficiently (Rybkin et al., 2019; Schmeckpeper et al., 2020; Sun et al., 2024). Consequently, humans can easily transfer observed actions to various contexts and imagine the transitions of new environments through a few interactions (Poggio & Bizzi, 2004). These insights motivate us to ponder: can we achieve human-like adaptability in world modeling by learning transferrable action representations from observations?
In this paper, we propose AdaWorld, an innovative pretraining approach for highly adaptable world models. Unlike previous methods that only pretrain on action-agnostic videos, we argue that incorporating action information during pretraining will significantly enhance the adaptability of world models. As illustrated in Figure 1, the adaptability of AdaWorld primarily manifests in two aspects: (1) Given one demonstration of an action, AdaWorld can readily transfer that action to various contexts without further training. (2) It can also be efficiently adapted into specialized world models through limited interactions and finetuning.
AdaWorld consists of two key components: a latent action autoencoder that extracts actions from unlabeled videos, and an autoregressive world model that takes the extracted actions as conditions. Our main challenge is that in-the-wild videos often involve complicated contexts (e.g., colors and textures), which hinders effective action recognition. To overcome this challenge, we introduce an information bottleneck design to our latent action autoencoder. Specifically, the latent action encoder extracts a compact encoding from two consecutive frames. We refer to this encoding as latent action hereinafter, as it is used to represent the transition between these two frames. Based on the latent action and the former frame, the latent action decoder makes its best effort to predict the subsequent frame. By minimizing the prediction loss using the minimal information encoded in the latent action, our autoencoder is encouraged to disentangle the most critical action from its context. Unlike previous methods (Bruce et al., 2024; Chen et al., 2024b; Ye et al., 2025) that focus on playability and behavior cloning, we compress the latent actions into a continuous latent space to maximize expressiveness and enable flexible composition. We find that our latent actions are context-invariant and can be effectively transferred across different contexts.
We then pretrain an autoregressive world model that conditions on the latent actions. Thanks to the strong transferability of our latent actions, the resulting world model is readily adaptable to various environments. In particular, since our world model has learned to simulate different actions represented by any latent actions, adapting to a new environment is akin to finding the corresponding latent actions for its action space. Given one demonstration of an action, our model can readily transfer the demonstrated action by extracting the latent action with latent action encoder and reusing it across different contexts. When action labels are provided, we can similarly derive their latent actions and use the averaged embedding to initialize specialized world models. This enables us to adapt our model to specialized world models with minimal finetuning. When only limited interactions are available (e.g., 50 interactions), our approach is significantly more efficient than pretraining from action-agnostic videos.
Note that our approach is as scalable as existing video pretraining methods (Seo et al., 2022; Wu et al., 2023; Agarwal et al., 2025; He et al., 2025; Yu et al., 2025). To enhance the generalization ability of AdaWorld, we collect a large corpus of videos from thousands of environments through automated generation. The resulting dataset encompasses extensive scenarios, spanning from ego perspectives and third-person views to virtual games and real-world activities. After action-aware pretraining at scale, we show that the adaptability of AdaWorld can seamlessly generalize to a wide variety of domains.
In summary, we make the following contributions:
  • We present AdaWorld, an autoregressive world model that is highly adaptable across various environments. It can readily transfer actions to different contexts and allows efficient adaptation with limited interactions.
  • We establish AdaWorld on a large-scale dataset sourced from extremely diverse environments. After extensive pretraining, AdaWorld demonstrates strong generalization capabilities across various domains.
  • We conduct comprehensive experiments across multiple environments to verify the efficacy of AdaWorld. Our model achieves promising results in action transfer, world model adaptation, and visual planning.
2 Method
In this section, we first introduce the architectural design of our latent action autoencoder (Sec. 2.1). By leveraging the latent actions as conditions, we then build an autoregressive world model through action-aware pretraining (Sec. 2.2). Finally, we demonstrate how our model facilitates highly adaptable world modeling (Sec. 2.3).
2.1 Latent Action Autoencoder
Instead of only taking actionless videos for world model pretraining, our key innovation is to incorporate action information during the pretraining phase. Benefiting from the pretrained knowledge of action controllability, the resulting world models can be efficiently adapted using limited ground truth actions. However, action labels are rarely available for in-the-wild videos. While a common practice is to collect these labels through interactions, collecting them across a multitude of environments incurs significant labor. Furthermore, defining a unified action format across diverse environments is often unfeasible, and current world models require costly training to accommodate new action formats.
Figure 2: Latent action autoencoder. With an information bottleneck design, our latent action autoencoder is able to extract the most critical action information from videos and compresses it into a continuous latent action.
To address these challenges, instead of relying on explicit action annotations, we propose extracting latent actions from videos as a unified condition for world model pretraining. Nevertheless, in general videos, the action information is often entangled with the contexts, posing significant difficulty for effective action recognition. Inspired by the observation that agents’ actions often drive the dominant variation in most interactive scenarios (Rybkin et al., 2019; Menapace et al., 2021, 2022; Bruce et al., 2024), we introduce an information bottleneck to automatically differentiate actions from observations.
To be specific, we instantiate a latent action autoencoder based on the Transformer architecture (Vaswani et al., 2017), where the encoder extracts the latent action $\tilde{a}$ from two consecutive frames $f_{t:t+1}$, and the decoder predicts the subsequent frame $f_{t+1}$ based on the latent action $\tilde{a}$ and the former frame $f_{t}$. The latent action encoder divides two frames $f_{t:t+1}$ into image patches of size $16\times 16$. These patches are then projected to patch embeddings and flattened along the spatial dimension. Afterwards, they are concatenated with two learnable tokens $a_{t:t+1}$. Sinusoidal position embeddings (Dosovitskiy et al., 2021) are also applied to each frame to indicate the spatial information. To efficiently encode the tokens from these two frames, we employ a spatiotemporal Transformer (Bruce et al., 2024) with $L$ stacked blocks. Each block comprises interleaved spatial and temporal attention modules, followed by a feed-forward network. The spatial attention can attend to all tokens within each frame, while the temporal attention has access to the two tokens in the same spatial positions across the two frames. We also incorporate rotary embeddings (Su et al., 2024) in temporal attentions to indicate the causal relationship. After sufficient attention correlations, the learnable tokens $a_{t:t+1}$ can adaptively aggregate the temporal dynamics between the two input frames. We take $a_{t+1}$ to predict the posterior of the latent action $(\mu_{\tilde{a}},\sigma_{\tilde{a}})$ following the standard VAE (Kingma & Welling, 2014). Subsequently, we sample $\tilde{a}$ from the approximated posterior and attach it to $f_{t}$, which is then sent to the latent action decoder. The latent action decoder is a spatial Transformer that predicts the subsequent frame $f_{t+1}$ in the pixel space. The whole latent action autoencoder is optimized with the VAE objective:
$\mathcal{L}^{pred}_{\theta,\phi}(f_{t+1})=$ $\;\mathbb{E}_{q_{\phi}(\tilde{a}|f_{t:t+1})}\log p_{\theta}(f_{t+1}|\tilde{a},f_{t})$ (1)
$-D_{KL}(q_{\phi}(\tilde{a}|f_{t:t+1})||p(\tilde{a})).$
Compared to the original pixel space, the dimension of our latent action is extremely compact. Hence, it is challenging to forward the entire subsequent frame to the decoder via latent action. To minimize the prediction loss of the subsequent frame, the latent action $\tilde{a}$ must encapsulate the most critical variations relative to the former frame. This results in context-invariant action representations that closely correspond to the true actions taken by the agents.
Figure 3: Action-aware pretraining. We extract latent actions from videos using the latent action encoder. By leveraging the extracted actions as unified conditions, we pretrain a world model that can perform autoregressive rollouts at inference.
Nevertheless, we empirically find that our latent action autoencoder, trained using the aforementioned formulation, struggles to express diverse transitions between frames. This problem arises because the standard VAE imposes a strong constraint on posterior distributions. Conversely, removing this constraint may compromise the disentanglement ability of VAE (Burgess et al., 2017). To remedy this, we adopt the $\beta$-VAE formulation (Higgins et al., 2017; Alemi et al., 2017), which introduces an adjustable hyperparameter $\beta$:
$\mathcal{L}^{pred}_{\theta,\phi}(f_{t+1})=$ $\;\mathbb{E}_{q_{\phi}(\tilde{a}|f_{t:t+1})}\log p_{\theta}(f_{t+1}|\tilde{a},f_{t})$ (2)
$-\beta\,D_{KL}(q_{\phi}(\tilde{a}|f_{t:t+1})||p(\tilde{a})).$
Figure 4: Action transfer from demonstrations. By extracting and reusing the latent actions in different contexts, AdaWorld can readily transfer the demonstrated actions from source videos to various target scenes without training. Please see Appendix C for more results.
The additional hyperparameter enables us to flexibly control the information contained by the latent actions. In practice, we adjust this hyperparameter to achieve a good trade-off between expressiveness and context disentangling ability of our latent actions. As shown in Figure 4, our latent action autoencoder can extract context-invariant actions that are transferrable across different contexts.
2.2 Action-Aware Pretraining
After training the latent action autoencoder, we can use its encoder to automatically extract action information from videos. This allows us to incorporate action information for world model pretraining, which we refer to as action-aware pretraining. To realize this, we pretrain a world model that predicts the next frame conditioned on the current latent action. As shown in Figure 3, we utilize the latent action encoder to extract the latent actions between frames and send them as inputs for our world model. Unlike previous methods that often predict video clips (Yang et al., 2024c; Xiang et al., 2024; Agarwal et al., 2025), our model supports frame-level control, offering finer granularity for interaction. To ensure smooth transitions, we maintain a short-term memory with $K$ historical frames. During inference, our model can predict a sequence of future frames by autoregressively repeating the next-frame prediction process and appending the predicted frames to the memory.
Figure 5: Action composition. We compose two observed actions by averaging their latent actions in the continuous latent space, resulting in a new action that combines the functionalities of both.
Although it may seem straightforward to repurpose the latent action decoder as the function of world model, it only makes coarse predictions via a single forward pass, resulting in significant quality degradation after several interactions. To achieve genuine predictions, we establish an independent world model based on diffusion models. Specifically, we initialize the world model with Stable Video Diffusion (SVD) (Blattmann et al., 2023), a latent diffusion model trained with the EDM framework (Karras et al., 2022). Different from the original SVD, we only denoise one noisy frame each time. To enable deep aggregation with the action information, the latent action is concatenated with both the timestep embedding and the CLIP image embedding from the original SVD. The last frame in memory is used as the condition image of SVD. To inherit the pretrained temporal modeling capability, we encode the historical frames using the SVD image encoder and concatenate them with the noise latent map of the frame to predict. Since the number of available historical frames may vary in practice, we randomly sample historical frames with a maximum length of 6 during training and provide the memory length to the world model. Following previous practices (He et al., 2022; Valevski et al., 2025), noise augmentation is also applied to corrupt the historical frames during training. This augmentation can effectively alleviate the long-term drift problem, even when no noise is applied during inference. We pretrain the world model on our large-scale dataset by minimizing the following diffusion loss:
$\mathcal{L}_{\text{pretrain}}=\mathbb{E}_{\boldsymbol{x}_{0},\epsilon,t}\Big{[}\|\boldsymbol{x}_{0}-\hat{\boldsymbol{x}}_{0}(\boldsymbol{x}_{t},t,\boldsymbol{c})\|^{2}\Big{]},$ (3)
where $\hat{\boldsymbol{x}}_{0}$ is the prediction of our world model and $\boldsymbol{c}$ is the conditioning information which includes historical frames and the latent action $\tilde{a}$.
2.3 High Adaptable World Models
After action-aware pretraining across various environments, the world model can be controlled by different latent actions, making it highly adaptable for multiple applications, including efficient action transfer and world model adaptation.
Efficient action transfer. When presented with a demonstration video, we use the latent action encoder to extract a sequence of latent actions. This enables us to disentangle the action from its context and replicate it across different contexts. Specifically, given the initial frame from a new context, we can reuse the extracted latent action sequence as the conditions to generate a new video autoregressively. As demonstrated in Figure 4, AdaWorld naturally transfers actions from source videos to various contexts.
Efficient world model adaptation. Our model also allows efficient world model adaptation with limited action labels and training steps. Specifically, after collecting a few action-video pairs through interactions, we use the latent action encoder to infer their latent actions. Thanks to the continuity of our latent action space, the action embedding for each ground truth action can be derived by simply averaging all corresponding latent actions. We empirically find that the averaged embedding consistently represents the intended action. Thus, for a new environment with $N$ action options, we initialize a specialized world model using $N$ averaged latent actions and finetune the whole model for a few steps. In Figure 6, we show that the models initialized in this way can be efficiently adapted through minimal finetuning.
Action composition and creation. It is also noteworthy that AdaWorld enables several unique applications compared to existing world models. For instance, it allows for the composition of new actions by interpolating observed actions within the latent space, as demonstrated in Figure 5. In addition, by collecting and clustering latent actions, we can easily create a flexible number of control options with distinct functions and strong controllability. This suggests that AdaWorld could serve as an alternative to generative interactive environments (Bruce et al., 2024). See Appendix C for experimental details on action creation.
3 Experiments
In this section, we first demonstrate AdaWorld’s strengths in action transfer in Sec. 3.1. We then study how efficient world model adaptation enables better simulation and planning in Sec. 3.2. Lastly, we analyze the effectiveness of our designs with ablation studies in Sec. 3.3.
To thoroughly understand the adaptability of our approach, we compare AdaWorld with three representative baselines:
  • Action-agnostic pretraining. In this setup, we pretrain a world model that shares the same architecture as AdaWorld but does not take latent actions as conditions. This baseline is used to demonstrate the effect of the predominant pretraining paradigm that relies only on action-agonistic videos (Wu et al., 2023; Gao et al., 2024; Che et al., 2025; Agarwal et al., 2025; He et al., 2025).
  • Optical flow as an action-aware condition. We automatically predict optical flows from videos using UniMatch (Xu et al., 2023a). The flow maps are downsampled to $16\times 16$ and flattened as conditional encodings to replace the latent actions during pretraining. This baseline serves as an alternative solution for extracting action information from unlabeled videos.
  • Discrete latent action as an action-aware condition. We also implement a variant of the latent action autoencoder based on the standard VQ-VAE (Van Den Oord et al., 2017). Instead of using a continuous latent action space, this variant adopts a VQ codebook with 8 discrete codes following Genie (Bruce et al., 2024).
Except for the above modifications, we align other training settings for the baselines and our method. The world models of all compared methods are trained for 50K iterations to ensure a fair comparison. Our training dataset comprises four publicly accessible datasets (Goyal et al., 2017; Grauman et al., 2022; O’Neill et al., 2024; Ju et al., 2024) and videos collected automatically from 1016 environments in Gym Retro (Nichol et al., 2018) and Procgen Benchmark (Cobbe et al., 2020). This results in approximately 2000 million frames of interactive scenarios in total. More details about our datasets and implementation are provided in Appendices A and B.
3.1 Action Transfer
AdaWorld can readily transfer a demonstrated action to various contexts without further training. Below, we provide both qualitative and quantitative evaluations to showcase how effectively AdaWorld performs action transfer.
Qualitative results. We transfer action sequences of length 20 through autoregressive generation in Figure 4. It shows that AdaWorld can effectively disentangle the demonstrated actions and emulate them across contexts. Qualitative comparison with other baselines can be found in Appendix C.
Quantitative results. To automatically compare the proposed method with other baselines, we construct a evaluation set sourced from the unseen LIBERO (Liu et al., 2023) and Something-Something v2 (SSv2) (Goyal et al., 2017) datasets. Specifically, we select and pair videos from the same tasks in LIBERO and the same labels among the top-10 most frequent labels in SSv2, resulting in 1300 pairs for evaluation (more details in Appendix D). While the selected video pairs contain similar actions, we find that the video pairs from LIBERO often differ in the arrangement of objects, and those from SSv2 have significant differences in contexts. For each video pair, we take the first video as the demonstration video and use the first frame from the second video as the initial frame. We then generate videos by extracting action conditions from the demonstration video and employing different models to autoregressively predict the next 20 frames from the initial frame. The evaluation is performed by measure the generated videos against the original videos using Fréchet Video Distance (FVD) (Unterthiner et al., 2018). To complement the FVD evaluation, which only reflects overall distribution similarity, we further employ Embedding Cosine Similarity (ECS) (Sun et al., 2024) that performs frame-level measurements based on the I3D features (Carreira & Zisserman, 2017). As reported in Table 1, our continuous latent action achieves the best action transfer performance, underscoring its capability to express more nuanced actions without losing generality.
Table 1: Action transfer comparison. In both datasets, AdaWorld excels at transferring the demonstrated actions to different contexts.
Method LIBERO SSv2
[…]
Act-agnostic 1545.2 0.702 847.2 0.592
Flow cond. 1409.5 0.724 702.8 0.611
Discrete cond. 1504.5 0.700 726.8 0.596
AdaWorld 767.0 0.804 473.4 0.639
Table 2: Adaptability comparison in three unseen environments. All compared models are finetuned by 800 steps using 100 samples for each action, where AdaWorld achieves the best performance with limited action-labeled data and finetuning steps.
Method Habitat Minecraft DMLab
[…]
Act-agnostic 20.34 0.450 19.44 0.532 20.96 0.386
Flow cond. 22.49 0.373 20.71 0.492 22.22 0.357
Discrete cond. 23.31 0.342 21.33 0.465 22.36 0.349
AdaWorld 23.58 0.327 21.59 0.457 22.92 0.335
3.2 World Model Adaptation
Figure 6: PSNR curves for world model adaptation in Minecraft. Despite learning from the same video sources, our AdaWorld adapts to new environments more efficiently than conventional pretraining methods.
We also investigate how the proposed method benefits efficient world model adaptation in terms of simulation quality and visual planning performance.
3.2.1 Simulation Quality
Setup. We choose three action-controlled environments (i.e., Habitat (Savva et al., 2019), Minecraft, DMLab (Beattie et al., 2016)) that are not included in our training dataset. To demonstrate the adaptability with restricted labels, we collect only 100 samples for each action in every environment. Using the limited interaction data, we then finetune all compared world models for 800 steps with a batch size of 32 and a learning rate of $5\times 10^{-5}$. The learning rate for the pretrained weights is discounted by a factor of 0.1. We randomly initialize action embeddings for the action-agnostic pretraining baseline. For the other three models, we use the averaged action conditions extracted from the 100 samples to initialize the action embeddings as described in Sec. 2.3. The adaptation quality is evaluated via PSNR (Hore & Ziou, 2010) and LPIPS (Zhang et al., 2018).
Results. As reported in Table 2, AdaWorld achieves the best fidelity after finetuning with limited interactions and compute. The comparison results suggest that the proposed method allows the world models to efficiently simulate new action controls in unseen environments. Note that all action-aware variants significantly outperform the action-agnostic baseline, underlining the importance of our key innovation, i.e., incorporating action information during pretraining.
To further demonstrate our sample efficiency and finetuning efficiency, we select Minecraft as the environment and conduct more comparative experiments using different sample numbers and finetuning steps. We show the evolving curves of PSNR in Figure 6. In all cases, AdaWorld performs much better at the beginning and improves significantly faster after a few finetuning steps. This suggests that our approach provides a superior initialization for efficient world model adaptation compared to conventional pretraining methods.
3.2.2 Visual Planning
Table 3: Visual planning results. For all selected scenes from four Procgen environments, we report the success rate and standard error with 5 random seeds. AdaWorld achieves higher average success rates than other baselines as well as Q-learning even without training. Oracle: We use the ground truth environments for MPC, which indicate the upper bound of this planning strategy.
[…]
Heist Jumper Maze CaveFlyer Average
Random 19.33$\pm$1.8022.00$\pm$1.0241.33$\pm$2.2222.00$\pm$1.0226.17$\pm$0.52%
Act-agnostic 20.67$\pm$1.4520.67$\pm$1.0039.33$\pm$1.1723.33$\pm$0.7526.00$\pm$0.20%
AdaWorld
   w/o finetune 38.67$\pm$0.8268.00$\pm$0.9241.33$\pm$1.1131.33$\pm$1.0244.83$\pm$0.28%
   w/ finetune 66.67$\pm$1.6758.67$\pm$1.0268.00$\pm$0.6933.33$\pm$1.5556.67$\pm$0.44%
Q-learning 22.67$\pm$1.5847.33$\pm$2.744.67$\pm$0.3334.00$\pm$2.5227.17$\pm$0.26%
Oracle (GT env.) 86.67$\pm$1.2977.33$\pm$1.0984.67$\pm$1.1974.00$\pm$1.6380.67$\pm$0.43%
Setup. After learning action controls, world models can be utilized for planning. To evaluate the impact of adaptability on planning performance, we compare AdaWorld with the common action-agnostic pretraining using a sampling-based model predictive control (MPC) optimized by Cross-Entropy Method (De Boer et al., 2005; Chua et al., 2018). The MPC planning and optimization procedure are deferred to Sec. B.4. We define a goal-reaching task based on the Procgen benchmark (Cobbe et al., 2020) and select 30 scenes from each of four environments (Heist, Jumper, Maze, CaveFlyer). This ensures that the specified goals can be reached within an acceptable number of steps (more details in Appendix E). For each scene, we randomly collect 100 samples for each action in the default action space (LEFT, DOWN, UP, RIGHT). Based on the collected samples, the pretrained world models are finetuned for 500 steps with a batch size of 32 and a learning rate of $5\times 10^{-5}$. We then use the finetuned world models to perform MPC planning in the selected scenes. The reward is defined as the cosine similarity between the predicted observations and the image of the final state. The MPC process is deemed successful if the agent reaches the final state within 20 steps.
Results. We present the success rates averaged over 5 random seeds in Table 3. While the action-agnostic baseline performs similarly to random planning, AdaWorld substantially increase the success rates across all environments. This indicates that our approach not only adapts more efficiently but also enhances visual planning performance.
Additionally, we evaluate the visual planning performance without finetuning our model using the collected samples. In particular, we only utilize the averaged latent actions derived from these samples as action embeddings for the corresponding scenes. The results in Table 3 indicate that even without updating model weights, our variant still outperforms the finetuned action-agnostic pretraining baseline.
To further demonstrate the effectiveness of our approach, we also compare the planning results with Q-learning (Sutton & Barto, 2018), a classical model-free reinforcement learning method. For each scene, we construct a Q-table using the same samples collected for the MPC planning. The states of Q-table are represented by quantized images, and the rewards are obtained by computing the cosine similarity with the goal image. As shown in Table 3, AdaWorld significantly dominates the Q-learning method, suggesting that our approach makes more effective use of limited interactions.
3.3 Ablation Study
Table 4: Impacts of training data diversity. Increasing data diversity enhances generalization to new domains.
Training Data Procgen
[…]
OpenX 25.51 0.318
Retro 26.43 0.250
Retro+OpenX 26.62 0.234
Averaged action embedding. We ablate the averaged action embedding proposed in 2.3 with random initialization and adapt the resultant model to the unseen Minecraft environment. From Figure 6, we observe that randomly initializing the action embeddings of AdaWorld results in a quality drop at the beginning. Nevertheless, it is noteworthy that although our variant is slightly worse than the action-agnostic pretraining baseline when the finetuning begins, it rapidly surpasses the action-agnostic baseline after just 200 steps. This is because AdaWorld has learned a highly adaptable control interface through action-aware pretraining, allowing it to efficiently adapt to an unseen environment by simply fitting the action embeddings for the new action space.
Data diversity. We also study the impact of data mixture to the generalization ability of latent actions. To this end, we implement three latent action autoencoder with different combinations of Open X-Embodiment (OpenX) (O’Neill et al., 2024) and Gym Retro (Nichol et al., 2018) datasets for 40K steps. We then assess the latent action decoder predictions of these three variants on Procgen benchmark (Cobbe et al., 2020) in Table 4. Surprisingly, we discover that even though OpenX mainly consists of real-world robot videos, incorporating OpenX helps the latent action autoencoder generalize to the unseen 2D virtual games in Procgen. This suggests that further increasing data diversity may positively affect the generalization of our latent actions.
4 Related Work
4.1 World Models
Driven by advances in deep generative models (Peebles & Xie, 2023; Yang et al., 2025), world models have made encouraging strides in recent years. They aim to replicate the transitions of the world through generation (Sutton, 1991; Ha & Schmidhuber, 2018; Yang et al., 2024c; Lu et al., 2024, 2025; Agarwal et al., 2025), enabling agents to perform decision making entirely in imagination even when confronted with unseen situations (Kaiser et al., 2020; Kim et al., 2020; Micheli et al., 2023; Hafner et al., 2023; Du et al., 2023; Mazzaglia et al., 2024). This capability has proven beneficial for gaming agents (Alonso et al., 2024; Pearce et al., 2024; He et al., 2025), autonomous vehicles (Hu et al., 2023; Yang et al., 2024a; Gao et al., 2024; Wang et al., 2024; Bar et al., 2025; Hassan et al., 2025), and real robots (Seo et al., 2023; Wu et al., 2022; Ko et al., 2024; Du et al., 2024; Zhen et al., 2024; Zhou et al., 2024b; Wu et al., 2024; Zhu et al., 2024; Zhou et al., 2024a; Bu et al., 2024; Wang et al., 2025; Qi et al., 2025; Wu et al., 2025).
Despite the substantial benefits, existing works often suffer from costly training that requires extensive annotations to learn new actions in different environments. To mitigate this issue, some efforts have explored more efficient world modeling through pretraining from passive videos (Watter et al., 2015; Seo et al., 2022; Mendonca et al., 2023; Wu et al., 2023). However, these methods primarily focus on obtaining compact world representations rather than learning control interfaces that can be readily adapted with limited interactions and computations. Another family of research investigates efficient online adaptation (Yang et al., 2024b; Rigter et al., 2024; Hong et al., 2025), but they primarily focus on parameter efficiency by assuming that the parameters of the pretrained video models are frozen or inaccessible.
While a recent study (Bruce et al., 2024) has explored learning interactive environments from 2D gameplay videos, it is limited to 8 fixed actions, which constrains its ability to express and adapt to more fine-grained actions. In contrast, we advocate for using a continuous latent action space, maximizing flexibility for efficient action transfer and enabling several unique applications for adaptable world modeling.
4.2 Latent Action from Videos
Humans can comprehend the notion of various actions by observing others to behave (Rizzolatti et al., 1996; Rybkin et al., 2019; Schmeckpeper et al., 2020). To develop intelligent agents, it is intriguing to automatically extract similar action primitives from videos. Some approaches align human and robot manipulation videos in a shared space to derive meaningful action prototypes (Xu et al., 2023b). However, they require collecting semantically paired videos, which limits their scalability for most real-world tasks.
To avoid reliance on video pairs, some works propose extracting implicit latent actions through fully unsupervised learning (Edwards et al., 2019; Menapace et al., 2021, 2022; Ye et al., 2023; Schmidt & Jiang, 2024; Sun et al., 2024; Villar-Corrales & Behnke, 2025). Nevertheless, they mainly focus on a single environment with limited complexity. While more recent works extend similar approaches to multiple datasets (Cui et al., 2024; Chen et al., 2024b, c; Ye et al., 2025; Bu et al., 2025; Ren et al., 2025), they mainly leverage latent actions as the objectives for behavior cloning. Consequently, they always adopt discrete actions to fit the policy networks (Kim et al., 2024), leading to significant ambiguity due to discretization (Nikulin et al., 2025). As a result, the potential of latent actions for adaptable world modeling remains underexplored.
5 Conclusion
In this paper, we introduce AdaWorld, a new world model learning approach that facilitates efficient adaptation across various environments. It is highly adaptable in transferring and learning new actions with limited interactions and finetuning. Extensive experiments and analyses demonstrate the superior adaptability of AdaWorld, highlighting its potential as a new paradigm for world model pretraining.
Limitations. While AdaWorld promotes adaptable world modeling, several challenges remain. First, it does not operate at real-time frequency. Future work could incorporate distillation and sampling techniques (Feng et al., 2024; Yin et al., 2025) to accelerate inference speed. Similar to prior works (Yang et al., 2024e), AdaWorld struggles to create novel content when the rollout exceeds the initial scene. This issue is likely to be solved by scaling model and training data (Bruce et al., 2024; Bar et al., 2025). Additionally, our model falls short in achieving extremely long-term rollouts, and we will explore potential solutions (Chen et al., 2024a; Feng et al., 2024; Ruhe et al., 2024) in future work. We also append some primary failure cases in Appendix C.
Impact Statement
This paper presents work whose goal is to advance the field of machine intelligence. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Datasets
A.1 Data Collection and Generation
Our training data is primarily sourced from four publicly accessible datasets (Goyal et al., 2017; Grauman et al., 2022; O’Neill et al., 2024; Ju et al., 2024) that represent various actions in typical interactive scenarios. For the videos from Open X-Embodiment (O’Neill et al., 2024), we keep both egocentric and exocentric demonstrations to encompass a wide spectrum of action patterns. Note that we manually remove the unfavorable subsets that are dominated by static, nonconsecutive, low-frequency and low-resolution frames.
To further enrich the diversity of our data, we also automatically generate a massive number of transitions from two gaming platforms developed by OpenAI (Nichol et al., 2018; Cobbe et al., 2020). We manually search on the Internet and import the ROMs we found into Gym Retro (Nichol et al., 2018), resulting in a total of 1000 interactive environments (see Table 6 for the full list). For the 16 environments in Procgen (Cobbe et al., 2020), we hold out 1000 start levels for evaluation and use the remaining 9000 levels for training. All samples are generated using the “hard” mode. We take a random action at each time step to collect 1M transitions for each Gym Retro environment and 10M for each Procgen environment. Unlike Genie (Bruce et al., 2024) that samples all actions uniformly in their case study, we employ a biased action sampling strategy to encourage broader exploration. Specifically, we increase the probability of selecting a particular action for a short period and then alternate these probabilities in the subsequent period. As shown in Figure 7, this simple strategy leads to much more diverse scenes in our data. Generating data with reinforcement learning agents may further boost the scene diversity (Kazemi et al., 2024; Yang et al., 2024d; Valevski et al., 2025), which we leave for future work. In Figure 8, we visualize some representative environments in our final dataset.
A.2 Data Mixture
Since the datasets vary in size and diversity, it is challenging to balance them perfectly. Therefore, we simply weight all subsets according to the number of videos during training. We report detailed statistics of our training data in Table 5.
Table 5: Data organization. Data sources, estimated frame counts, and mixture ratios for our training.
Category Data Source # Frames Ratios
Robot Data Open X-Embodiment (O’Neill et al., 2024) 170M 30%
Human Activity Ego4D (Grauman et al., 2022) 330M 1%
Something-Something V2 (Goyal et al., 2017) 7M 3%
2D Video Game Gym Retro (Nichol et al., 2018) 1000M 49%
Procgen Benchmark (Cobbe et al., 2020) 144M 2%
3D Rendering MiraData (Ju et al., 2024) 200M 14%
City Walking MiraData (Ju et al., 2024) 120M 1%
Figure 7: Effect of our biased action sampling strategy. Compared to the uniform action sampling strategy, our biased scheme enables agents to explore longer horizons.
Table 6: Full list of collected Gym Retro environments. We import 1000 ROMs from the web for automated data generation at scale.
1942-Nes 1943-Nes 3NinjasKickBack-Genesis 8Eyes-Nes AaahhRealMonsters-Genesis AbadoxTheDeadlyInnerWar-Nes AcceleBrid-Snes
ActRaiser2-Snes ActionPachio-Snes AddamsFamily-GameBoy AddamsFamily-Genesis AddamsFamily-Nes AddamsFamily-Sms AddamsFamily-Snes
AddamsFamilyPugsleysScavengerHunt-Nes AddamsFamilyPugsleysScavengerHunt-Snes AdvancedBusterhawkGleylancer-Genesis Adventure-Atari2600 AdventureIsland-GameBoy AdventureIsland3-Nes AdventureIslandII-Nes
AdventuresOfBatmanAndRobin-Genesis AdventuresOfBayouBilly-Nes AdventuresOfDinoRiki-Nes AdventuresOfDrFranken-Snes AdventuresOfKidKleets-Snes AdventuresOfMightyMax-Genesis AdventuresOfMightyMax-Snes
AdventuresOfRockyAndBullwinkleAndFriends-Genesis AdventuresOfRockyAndBullwinkleAndFriends-Nes AdventuresOfRockyAndBullwinkleAndFriends-Snes AdventuresOfStarSaver-GameBoy AdventuresOfYogiBear-Snes AeroFighters-Snes AeroStar-GameBoy
AeroTheAcroBat-Snes AeroTheAcroBat2-Genesis AeroTheAcroBat2-Snes AfterBurnerII-Genesis AfterBurst-GameBoy AirBuster-Genesis AirCavalry-Snes
AirDiver-Genesis AirRaid-Atari2600 Airstriker-Genesis Airwolf-Nes AlfredChicken-GameBoy AlfredChicken-Nes AlfredChicken-Snes
Alien-Atari2600 Alien3-Nes Alien3-Sms AlienSoldier-Genesis AlienSyndrome-Sms AlienVsPredator-Snes Alleyway-GameBoy
AlphaMission-Nes AlteredBeast-Genesis Amagon-Nes AmazingPenguin-GameBoy AmericanGladiators-Genesis Amidar-Atari2600 ArchRivalsTheArcadeGame-Genesis
ArcherMacleansSuperDropzone-Snes ArdyLightfoot-Snes Argus-Nes ArielTheLittleMermaid-Genesis Arkanoid-Nes ArkistasRing-Nes Armadillo-Nes
ArrowFlash-Genesis ArtOfFighting-Genesis ArtOfFighting-Snes Assault-Atari2600 Asterix-Atari2600 Asterix-Sms Asterix-Snes
AsterixAndObelix-GameBoy AsterixAndTheGreatRescue-Genesis AsterixAndTheGreatRescue-Sms AsterixAndThePowerOfTheGods-Genesis AsterixAndTheSecretMission-Sms Asteroids-Atari2600 Asteroids-GameBoy
AstroRabby-GameBoy AstroRoboSasa-Nes AstroWarrior-Sms Astyanax-Nes Athena-Nes Atlantis-Atari2600 AtlantisNoNazo-Nes
AtomicRoboKid-Genesis AtomicRunner-Genesis AttackAnimalGakuen-Nes AttackOfTheKillerTomatoes-GameBoy AttackOfTheKillerTomatoes-Nes AwesomePossumKicksDrMachinosButt-Genesis Axelay-Snes
AyrtonSennasSuperMonacoGPII-Genesis BOB-Genesis BOB-Snes BWings-Nes BackToTheFuturePartIII-Genesis BadDudes-Nes BadStreetBrawler-Nes
BakuretsuSenshiWarrior-GameBoy BalloonFight-Nes BalloonKid-GameBoy Baltron-Nes BananaPrince-Nes BanishingRacer-GameBoy BankHeist-Atari2600
Barbie-Nes BarkleyShutUpAndJam-Genesis BarkleyShutUpAndJam2-Genesis BarneysHideAndSeekGame-Genesis BartSimpsonsEscapeFromCampDeadly-GameBoy Batman-Genesis BatmanReturns-Genesis
BatmanReturns-Nes BatmanReturns-Snes BattleArenaToshinden-GameBoy BattleBull-GameBoy BattleCity-Nes BattleMasterKyuukyokuNoSenshiTachi-Snes BattleSquadron-Genesis
BattleTechAGameOfArmoredCombat-Genesis BattleUnitZeoth-GameBoy BattleZequeDen-Snes BattleZone-Atari2600 Battletoads-Genesis Battletoads-Nes BattletoadsDoubleDragon-Genesis
BattletoadsDoubleDragon-Snes BattletoadsInBattlemaniacs-Snes BattletoadsInRagnaroksWorld-GameBoy BeamRider-Atari2600 BeautyAndTheBeastBellesQuest-Genesis BeautyAndTheBeastRoarOfTheBeast-Genesis BebesKids-Snes
Berzerk-Atari2600 BillAndTedsExcellentGameBoyAdventure-GameBoy BiminiRun-Genesis BinaryLand-Nes BioHazardBattle-Genesis BioMetal-Snes BioMiracleBokutteUpa-Nes
BioSenshiDanIncreaserTonoTatakai-Nes BirdWeek-Nes BishoujoSenshiSailorMoon-Genesis BishoujoSenshiSailorMoonR-Snes BlaZeonTheBioCyborgChallenge-Snes BladeEagle-Sms BladesOfVengeance-Genesis
BlockKuzushi-Snes BlockKuzushiGB-GameBoy Blockout-Genesis BodyCount-Genesis BombJack-GameBoy BomberRaid-Sms BonkersWaxUp-Sms
BoobyBoys-GameBoy BoobyKids-Nes BoogermanAPickAndFlickAdventure-Genesis BoogermanAPickAndFlickAdventure-Snes BoogieWoogieBowling-Genesis BoulderDash-GameBoy BoulderDash-Nes
Bowling-Atari2600 Boxing-Atari2600 BoxingLegendsOfTheRing-Genesis BramStokersDracula-Genesis BramStokersDracula-Nes BramStokersDracula-Snes BrawlBrothers-Snes
BreakThru-Nes Breakout-Atari2600 BronkieTheBronchiasaurus-Snes BubbaNStix-Genesis BubbleAndSqueak-Genesis BubbleBobble-Nes BubbleBobble-Sms
BubbleBobblePart2-Nes BubbleGhost-GameBoy BubsyII-Genesis BubsyII-Snes BubsyInClawsEncountersOfTheFurredKind-Genesis BubsyInClawsEncountersOfTheFurredKind-Snes BuckyOHare-Nes
BugsBunnyBirthdayBlowout-Nes BullsVersusBlazersAndTheNBAPlayoffs-Genesis BullsVsLakersAndTheNBAPlayoffs-Genesis BuraiFighter-Nes BurningForce-Genesis CacomaKnightInBizyland-Snes Cadash-Genesis
CalRipkenJrBaseball-Genesis Caliber50-Genesis CaliforniaGames-Genesis Cameltry-Snes CannonFodder-Genesis CannonFodder-Snes CaptainAmericaAndTheAvengers-Genesis
CaptainAmericaAndTheAvengers-Nes CaptainAmericaAndTheAvengers-Snes CaptainCommando-Snes CaptainPlanetAndThePlaneteers-Genesis CaptainPlanetAndThePlaneteers-Nes CaptainSilver-Nes Carnival-Atari2600
Castelian-Nes CastleOfIllusion-Genesis Castlevania-Nes CastlevaniaBloodlines-Genesis CastlevaniaDraculaX-Snes CastlevaniaIIIDraculasCurse-Nes CastlevaniaTheNewGeneration-Genesis
CatNindenTeyandee-Nes Centipede-Atari2600 ChacknPop-Nes Challenger-Nes ChampionsWorldClassSoccer-Genesis ChampionshipProAm-Genesis ChaosEngine-Genesis
ChaseHQII-Genesis CheeseCatAstropheStarringSpeedyGonzales-Genesis CheeseCatAstropheStarringSpeedyGonzales-Sms ChesterCheetahTooCoolToFool-Genesis ChesterCheetahTooCoolToFool-Snes ChesterCheetahWildWildQuest-Genesis ChesterCheetahWildWildQuest-Snes
ChiChisProChallengeGolf-Genesis ChikiChikiBoys-Genesis Choplifter-Nes ChoplifterIIIRescueSurvive-Snes ChopperCommand-Atari2600 ChouFuyuuYousaiExedExes-Nes ChoujikuuYousaiMacross-Nes
ChoujikuuYousaiMacrossScrambledValkyrie-Snes ChubbyCherub-Nes ChuckRock-Genesis ChuckRock-Sms ChuckRock-Snes ChuckRockIISonOfChuck-Genesis ChuckRockIISonOfChuck-Sms
CircusCaper-Nes CircusCharlie-Nes CityConnection-Nes ClayFighter-Genesis Claymates-Snes Cliffhanger-Genesis Cliffhanger-Nes
Cliffhanger-Snes CloudMaster-Sms CluCluLand-Nes CobraTriangle-Nes CodeNameViper-Nes CollegeSlam-Genesis Columns-Genesis
ColumnsIII-Genesis CombatCars-Genesis ComicalMachineGunJoe-Sms ComixZone-Genesis Conan-Nes CongosCaper-Snes ConquestOfTheCrystalPalace-Nes
ContraForce-Nes CoolSpot-Genesis CoolSpot-Sms CoolSpot-Snes CosmicEpsilon-Nes CosmoGangTheVideo-Snes CrackDown-Genesis
CrazyClimber-Atari2600 CrossFire-Nes CrueBallHeavyMetalPinball-Genesis Curse-Genesis CutieSuzukiNoRingsideAngel-Genesis CutthroatIsland-Genesis CyberShinobi-Sms
Cyberball-Genesis Cybernator-Snes CyborgJustice-Genesis DJBoy-Genesis DaffyDuckInHollywood-Sms DaffyDuckTheMarvinMissions-Snes DaisenpuuTwinHawk-Genesis
DananTheJungleFighter-Sms DangerousSeed-Genesis DariusForce-Snes DariusII-Genesis DariusTwin-Snes DarkCastle-Genesis Darkman-Nes
Darwin4081-Genesis DashGalaxyInTheAlienAsylum-Nes DashinDesperadoes-Genesis DavidRobinsonsSupremeCourt-Genesis DazeBeforeChristmas-Genesis DazeBeforeChristmas-Snes DeadlyMoves-Genesis
DeathDuel-Genesis DeepDuckTroubleStarringDonaldDuck-Sms Defender-Atari2600 DefenderII-Nes DemonAttack-Atari2600 DennisTheMenace-Snes DesertStrikeReturnToTheGulf-Genesis
DevilCrashMD-Genesis DevilishTheNextPossession-Genesis DickTracy-Genesis DickTracy-Sms DickVitalesAwesomeBabyCollegeHoops-Genesis DigDugIITroubleInParadise-Nes DiggerTheLegendOfTheLostCity-Nes
DimensionForce-Snes DinoCity-Snes DinoLand-Genesis DirtyHarry-Nes DonDokoDon-Nes DonkeyKong-Nes DonkeyKong3-Nes
DonkeyKongCountry-Snes DonkeyKongCountry2-Snes DonkeyKongCountry3DixieKongsDoubleTrouble-Snes DonkeyKongJr-Nes DoubleDragon-Genesis DoubleDragon-Nes DoubleDragonIITheRevenge-Genesis
DoubleDragonIITheRevenge-Nes DoubleDragonVTheShadowFalls-Genesis DoubleDribbleThePlayoffEdition-Genesis DoubleDunk-Atari2600 DrRobotniksMeanBeanMachine-Genesis DragonPower-Nes DragonSpiritTheNewLegend-Nes
DragonTheBruceLeeStory-Genesis DragonTheBruceLeeStory-Snes DragonsLair-Snes DragonsRevenge-Genesis DreamTeamUSA-Genesis DynamiteDuke-Genesis DynamiteDuke-Sms
DynamiteHeaddy-Genesis ESPNBaseballTonight-Genesis EarnestEvans-Genesis EarthDefenseForce-Snes ElViento-Genesis ElementalMaster-Genesis ElevatorAction-Atari2600
ElevatorAction-Nes EliminateDown-Genesis Enduro-Atari2600 EuropeanClubSoccer-Genesis ExMutants-Genesis Exerion-Nes F1-Genesis
FZSenkiAxisFinalZone-Genesis FaeryTaleAdventure-Genesis FamilyDog-Snes Fantasia-Genesis FantasticDizzy-Genesis FantasyZone-Sms FantasyZoneIIOpaOpaNoNamida-Nes
FantasyZoneIITheTearsOfOpaOpa-Sms FantasyZoneTheMaze-Sms FatalFury-Genesis FatalFury2-Genesis FatalLabyrinth-Genesis FatalRewind-Genesis FelixTheCat-Nes
FerrariGrandPrixChallenge-Genesis FightingMasters-Genesis FinalBubbleBobble-Sms FinalFight-Snes FinalFight2-Snes FinalFight3-Snes FinalFightGuy-Snes
FireAndIce-Sms FirstSamurai-Snes FishingDerby-Atari2600 FistOfTheNorthStar-Nes Flicky-Genesis Flintstones-Genesis Flintstones-Snes
FlintstonesTheRescueOfDinoAndHoppy-Nes FlyingDragonTheSecretScroll-Nes FlyingHero-Nes FlyingHeroBugyuruNoDaibouken-Snes ForemanForReal-Genesis ForgottenWorlds-Genesis FormationZ-Nes
FoxsPeterPanAndThePiratesTheRevengeOfCaptainHook-Nes FrankThomasBigHurtBaseball-Genesis Freeway-Atari2600 Frogger-Genesis FrontLine-Nes Frostbite-Atari2600 FushigiNoOshiroPitPot-Sms
GIJoeARealAmericanHero-Nes GIJoeTheAtlantisFactor-Nes GadgetTwins-Genesis Gaiares-Genesis GainGround-Genesis GalagaDemonsOfDeath-Nes GalaxyForce-Sms
GalaxyForceII-Genesis Gauntlet-Genesis Gauntlet-Sms Geimos-Nes GeneralChaos-Genesis GhostsnGoblins-Nes GhoulSchool-Nes
GhoulsnGhosts-Genesis Gimmick-Nes GlobalDefense-Sms GlobalGladiators-Genesis Gods-Genesis Gods-Snes GokujouParodius-Snes
GoldenAxe-Genesis GoldenAxeIII-Genesis Gopher-Atari2600 Gradius-Nes GradiusII-Nes GradiusIII-Snes GradiusTheInterstellarAssault-GameBoy
Granada-Genesis Gravitar-Atari2600 GreatCircusMysteryStarringMickeyAndMinnie-Genesis GreatCircusMysteryStarringMickeyAndMinnie-Snes GreatTank-Nes GreatWaldoSearch-Genesis GreendogTheBeachedSurferDude-Genesis
Gremlins2TheNewBatch-Nes GrindStormer-Genesis Growl-Genesis GuardianLegend-Nes GuerrillaWar-Nes GunNac-Nes Gunship-Genesis
Gynoug-Genesis Gyrodine-Nes Gyruss-Nes HammerinHarry-Nes HangOn-Sms HardDrivin-Genesis HarleysHumongousAdventure-Snes
Havoc-Genesis HeavyBarrel-Nes HeavyNova-Genesis HeavyUnitMegaDriveSpecial-Genesis Hellfire-Genesis HelloKittyWorld-Nes Hero-Atari2600
HighStakesGambling-GameBoy HomeAlone-Genesis HomeAlone2LostInNewYork-Genesis HomeAlone2LostInNewYork-Nes Hook-Genesis Hook-Snes HuntForRedOctober-Nes
HuntForRedOctober-Snes Hurricanes-Genesis Hurricanes-Snes IMGInternationalTourTennis-Genesis IceClimber-Nes IceHockey-Atari2600 Ikari-Nes
IkariIIITheRescue-Nes Ikki-Nes Imperium-Snes Incantation-Snes IncredibleCrashDummies-Genesis IncredibleHulk-Genesis IncredibleHulk-Sms
IncredibleHulk-Snes IndianaJonesAndTheLastCrusade-Genesis IndianaJonesAndTheTempleOfDoom-Nes InsectorX-Genesis InsectorX-Nes InspectorGadget-Snes IronSwordWizardsAndWarriorsII-Nes
IshidoTheWayOfStones-Genesis IsolatedWarrior-Nes ItchyAndScratchyGame-Snes IzzysQuestForTheOlympicRings-Genesis IzzysQuestForTheOlympicRings-Snes Jackal-Nes JackieChansActionKungFu-Nes
JajamaruNoDaibouken-Nes JamesBond007TheDuel-Genesis JamesBond007TheDuel-Sms JamesBondJr-Nes JamesPond2CodenameRoboCod-Sms JamesPond3-Genesis JamesPondIICodenameRobocod-Genesis
JamesPondUnderwaterAgent-Genesis Jamesbond-Atari2600 Jaws-Nes JellyBoy-Snes JetsonsCogswellsCaper-Nes JetsonsInvasionOfThePlanetPirates-Snes JewelMaster-Genesis
JoeAndMac-Genesis JoeAndMac-Nes JoeAndMac-Snes JoeAndMac2LostInTheTropics-Snes JourneyEscape-Atari2600 JourneyToSilius-Nes Joust-Nes
JuJuDensetsuTokiGoingApeSpit-Genesis JudgeDredd-Genesis JudgeDredd-Snes JungleBook-Genesis JungleBook-Nes JungleBook-Snes JusticeLeagueTaskForce-Genesis
KaGeKiFistsOfSteel-Genesis Kaboom-Atari2600 KabukiQuantumFighter-Nes KaiketsuYanchaMaru2KarakuriLand-Nes KaiketsuYanchaMaru3TaiketsuZouringen-Nes KaiteTsukutteAsoberuDezaemon-Snes KamenNoNinjaAkakage-Nes
Kangaroo-Atari2600 KanshakudamaNageKantarouNoToukaidouGojuusanTsugi-Nes KeroppiToKeroriinuNoSplashBomb-Nes KidChameleon-Genesis KidIcarus-Nes KidKlownInCrazyChase-Snes KidKlownInNightMayorWorld-Nes
KidNikiRadicalNinja-Nes KingOfDragons-Snes KingOfTheMonsters2-Genesis KingOfTheMonsters2-Snes KirbysAdventure-Nes KnightsOfTheRound-Snes Krull-Atari2600
KungFu-Nes KungFuHeroes-Nes KungFuKid-Sms KungFuMaster-Atari2600 LandOfIllusionStarringMickeyMouse-Sms LastActionHero-Genesis LastActionHero-Nes
LastActionHero-Snes LastBattle-Genesis LastStarfighter-Nes LawnmowerMan-Genesis Legend-Snes LegendOfGalahad-Genesis LegendOfKage-Nes
LegendOfPrinceValiant-Nes LegendaryWings-Nes LethalEnforcers-Genesis LethalEnforcersIIGunFighters-Genesis LethalWeapon-Snes LifeForce-Nes LineOfFire-Sms
LittleMermaid-Nes LowGManTheLowGravityMan-Nes LuckyDimeCaperStarringDonaldDuck-Sms MCKids-Nes MUSHA-Genesis MagicBoy-Snes MagicSword-Snes
MagicalQuestStarringMickeyMouse-Snes MagicalTaruruutoKun-Genesis Magmax-Nes MaouRenjishi-Genesis MappyLand-Nes MarbleMadness-Genesis MarbleMadness-Sms
MarioBros-Nes Marko-Genesis Marsupilami-Genesis MarvelLand-Genesis Mask-Snes MasterOfDarkness-Sms McDonaldsTreasureLandAdventure-Genesis
MechanizedAttack-Nes MegaMan-Nes MegaMan2-Nes MegaManTheWilyWars-Genesis MegaSWIV-Genesis MegaTurrican-Genesis MendelPalace-Nes
MetalStorm-Nes MichaelJacksonsMoonwalker-Genesis MichaelJacksonsMoonwalker-Sms MickeyMousecapade-Nes MidnightResistance-Genesis MightyBombJack-Nes MightyFinalFight-Nes
MightyMorphinPowerRangers-Genesis MightyMorphinPowerRangersTheMovie-Genesis MightyMorphinPowerRangersTheMovie-Snes Millipede-Nes MitsumeGaTooru-Nes MoeroTwinBeeCinnamonHakaseOSukue-Nes MonsterInMyPocket-Nes
MonsterLair-Genesis MonsterParty-Nes MontezumaRevenge-Atari2600 MoonPatrol-Atari2600 MortalKombat-Genesis MortalKombat3-Genesis MortalKombatII-Genesis
MrNutz-Genesis MrNutz-Snes MsPacMan-Genesis MsPacMan-Nes MsPacMan-Sms MsPacman-Atari2600 MutantVirusCrisisInAComputerWorld-Nes
MyHero-Sms MysteryQuest-Nes NARC-Nes NHL94-Genesis NHL941on1-Genesis NameThisGame-Atari2600 NewZealandStory-Genesis
NewZealandStory-Sms Ninja-Sms NinjaCrusaders-Nes NinjaGaiden-Nes NinjaGaiden-Sms NinjaGaidenIIITheAncientShipOfDoom-Nes NinjaGaidenIITheDarkSwordOfChaos-Nes
NinjaKid-Nes NoahsArk-Nes NormysBeachBabeORama-Genesis OperationWolf-Nes Ottifants-Genesis Ottifants-Sms OutToLunch-Snes
OverHorizon-Nes POWPrisonersOfWar-Nes PacInTime-Snes PacManNamco-Nes PacMania-Genesis PacMania-Sms PanicRestaurant-Nes
Paperboy-Genesis Paperboy-Nes Paperboy-Sms Paperboy2-Genesis Parodius-Nes Parodius-Snes PeaceKeepers-Snes
PenguinKunWars-Nes Phalanx-Snes Phelios-Genesis Phoenix-Atari2600 PinkGoesToHollywood-Genesis PiratesOfDarkWater-Snes PitFighter-Genesis
PitFighter-Sms Pitfall-Atari2600 PitfallTheMayanAdventure-Genesis PitfallTheMayanAdventure-Snes PizzaPop-Nes Plok-Snes Pong-Atari2600
Pooyan-Atari2600 Pooyan-Nes Popeye-Nes PopnTwinBee-Snes PopnTwinBeeRainbowBellAdventures-Snes PorkyPigsHauntedHoliday-Snes PoseidonWars3D-Sms
PowerAthlete-Genesis PowerPiggsOfTheDarkAge-Snes PowerStrike-Sms PowerStrikeII-Sms Predator2-Genesis Predator2-Sms PrehistorikMan-Snes
PrivateEye-Atari2600 PsychicWorld-Sms Pulseman-Genesis PunchOut-Nes Punisher-Genesis Punisher-Nes PussNBootsPerosGreatAdventure-Nes
PuttySquad-Snes QBert-Nes Qbert-Atari2600 QuackShot-Genesis Quartet-Sms Quarth-Nes RType-Sms
RTypeIII-Snes RadicalRex-Genesis RadicalRex-Snes RaidenDensetsu-Snes RaidenDensetsuRaidenTrad-Genesis RainbowIslands-Nes RainbowIslandsStoryOfTheBubbleBobble2-Sms
RamboFirstBloodPartII-Sms RamboIII-Genesis Rampage-Nes RangerX-Genesis RastanSagaII-Genesis Realm-Snes RenAndStimpyShowPresentsStimpysInvention-Genesis
RenAndStimpyShowVeediots-Snes RenderingRangerR2-Snes Renegade-Nes Renegade-Sms RevengeOfShinobi-Genesis RiseOfTheRobots-Genesis RiskyWoods-Genesis
Ristar-Genesis RivalTurf-Snes Riverraid-Atari2600 RoadRunner-Atari2600 RoadRunnersDeathValleyRally-Snes RoboCop2-Nes RoboCop3-Genesis
RoboCop3-Nes RoboCop3-Sms RoboCop3-Snes RoboCopVersusTheTerminator-Sms RoboCopVersusTheTerminator-Snes RoboWarrior-Nes RoboccoWars-Nes
Robotank-Atari2600 RocketKnightAdventures-Genesis RockinKats-Nes Rollerball-Nes Rollergames-Nes RollingThunder2-Genesis RunSaber-Snes
RunningBattle-Sms RushnAttack-Nes SCATSpecialCyberneticAttackTeam-Nes SDHeroSoukessenTaoseAkuNoGundan-Nes Sagaia-Genesis Sagaia-Sms SaintSword-Genesis
SameSameSame-Genesis SamuraiShodown-Genesis Sansuu5And6NenKeisanGame-Nes Satellite7-Sms Seaquest-Atari2600 SecondSamurai-Genesis SectionZ-Nes
Seicross-Nes SeikimaIIAkumaNoGyakushuu-Nes SenjouNoOokamiIIMercs-Genesis ShadowDancerTheSecretOfShinobi-Genesis ShadowOfTheBeast-Genesis ShaqFu-Genesis Shatterhand-Nes
Shinobi-Sms ShinobiIIIReturnOfTheNinjaMaster-Genesis SilverSurfer-Nes SimpsonsBartVsTheSpaceMutants-Genesis SimpsonsBartVsTheSpaceMutants-Nes SimpsonsBartVsTheWorld-Nes SimpsonsBartmanMeetsRadioactiveMan-Nes
SkeletonKrew-Genesis Skiing-Atari2600 SkuljaggerRevoltOfTheWesticans-Snes SkyDestroyer-Nes SkyKid-Nes SkyShark-Nes SmartBall-Snes
SmashTV-Nes Smurfs-Genesis Smurfs-Nes Smurfs-Snes SnakeRattleNRoll-Nes SnowBrothers-Nes Socket-Genesis
SolDeace-Genesis Solaris-Atari2600 SoldiersOfFortune-Genesis SonSon-Nes SonicAndKnuckles-Genesis SonicAndKnuckles3-Genesis SonicBlast-Sms
SonicBlastMan-Snes SonicBlastManII-Snes SonicTheHedgehog-Genesis SonicTheHedgehog-Sms SonicTheHedgehog2-Genesis SonicTheHedgehog2-Sms SonicTheHedgehog3-Genesis
SonicWings-Snes SpaceHarrier-Nes SpaceHarrier-Sms SpaceHarrier3D-Sms SpaceHarrierII-Genesis SpaceInvaders-Atari2600 SpaceInvaders-Nes
SpaceInvaders-Snes SpaceInvaders91-Genesis SpaceMegaforce-Snes SpankysQuest-Snes Sparkster-Genesis Sparkster-Snes SpartanX2-Nes
SpeedyGonzalesLosGatosBandidos-Snes Spelunker-Nes SpiderManReturnOfTheSinisterSix-Sms Splatterhouse2-Genesis SpotGoesToHollywood-Genesis SprigganPowered-Snes SpyHunter-Nes
Sqoon-Nes StarForce-Nes StarGunner-Atari2600 StarSoldier-Nes StarWars-Nes StarshipHector-Nes SteelEmpire-Genesis
SteelTalons-Genesis Stinger-Nes StoneProtectors-Snes Stormlord-Genesis StreetFighterIISpecialChampionEdition-Genesis StreetSmart-Genesis StreetsOfRage-Genesis
StreetsOfRage2-Genesis StreetsOfRage3-Genesis StreetsOfRageII-Sms Strider-Genesis SubTerrania-Genesis SubmarineAttack-Sms SunsetRiders-Genesis
SuperAdventureIsland-Snes SuperAlfredChicken-Snes SuperArabian-Nes SuperBCKid-Snes SuperC-Nes SuperCastlevaniaIV-Snes SuperDoubleDragon-Snes
SuperFantasyZone-Genesis SuperGhoulsnGhosts-Snes SuperHangOn-Genesis SuperJamesPond-Snes SuperMarioBros-Nes SuperMarioBros2Japan-Nes SuperMarioBros3-Nes
SuperMarioWorld-Snes SuperMarioWorld2-Snes SuperPitfall-Nes SuperRType-Snes SuperSWIV-Snes SuperSmashTV-Genesis SuperSmashTV-Snes
SuperSpaceInvaders-Sms SuperStarForce-Nes SuperStarWars-Snes SuperStarWarsReturnOfTheJedi-Snes SuperStarWarsTheEmpireStrikesBack-Snes SuperStrikeGunner-Snes SuperThunderBlade-Genesis
SuperTrollIslands-Snes SuperTurrican-Snes SuperTurrican2-Snes SuperValisIV-Snes SuperWidget-Snes SuperWonderBoy-Sms SuperXeviousGumpNoNazo-Nes
Superman-Genesis SupermanTheManOfSteel-Sms SwampThing-Nes SwordOfSodan-Genesis SydOfValis-Genesis SylvesterAndTweetyInCageyCapers-Genesis T2TheArcadeGame-Genesis
T2TheArcadeGame-Sms TaiyouNoYuushaFighbird-Nes TakahashiMeijinNoBugutteHoney-Nes TargetEarth-Genesis TargetRenegade-Nes TaskForceHarrierEX-Genesis TazMania-Genesis
TazMania-Sms TazMania-Snes TeenageMutantNinjaTurtles-Nes TeenageMutantNinjaTurtlesIIITheManhattanProject-Nes TeenageMutantNinjaTurtlesIITheArcadeGame-Nes TeenageMutantNinjaTurtlesIVTurtlesInTime-Snes TeenageMutantNinjaTurtlesTheHyperstoneHeist-Genesis
TeenageMutantNinjaTurtlesTournamentFighters-Genesis TeenageMutantNinjaTurtlesTournamentFighters-Nes Tennis-Atari2600 Terminator-Genesis Terminator-Sms Terminator2JudgmentDay-Nes TerraCresta-Nes
TetrastarTheFighter-Nes Tetris-GameBoy TetrisAttack-Snes TetsuwanAtom-Nes Thexder-Nes ThunderAndLightning-Nes ThunderBlade-Sms
ThunderForceII-Genesis ThunderForceIII-Genesis ThunderForceIV-Genesis ThunderFox-Genesis ThunderSpirits-Snes Thundercade-Nes Tick-Genesis
Tick-Snes TigerHeli-Nes TimePilot-Atari2600 TimeZone-Nes TinHead-Genesis TinyToonAdventures-Nes TinyToonAdventuresBusterBustsLoose-Snes
TinyToonAdventuresBustersHiddenTreasure-Genesis Toki-Nes TomAndJerry-Snes TotalRecall-Nes TotallyRad-Nes ToxicCrusaders-Genesis ToxicCrusaders-Nes
TrampolineTerror-Genesis TransBot-Sms TreasureMaster-Nes Trog-Nes Trojan-Nes TrollsInCrazyland-Nes TroubleShooter-Genesis
Truxton-Genesis Turrican-Genesis Tutankham-Atari2600 TwinBee-Nes TwinBee3PokoPokoDaimaou-Nes TwinCobra-Nes TwinCobraDesertAttackHelicopter-Genesis
TwinEagle-Nes TwinkleTale-Genesis TwoCrudeDudes-Genesis UNSquadron-Snes UchuuNoKishiTekkamanBlade-Snes UndeadLine-Genesis UniversalSoldier-Genesis
Untouchables-Nes UpNDown-Atari2600 UrbanChampion-Nes UruseiYatsuraLumNoWeddingBell-Nes UzuKeobukseon-Genesis VRTroopers-Genesis VaporTrail-Genesis
Vectorman-Genesis Vectorman2-Genesis Venture-Atari2600 ViceProjectDoom-Nes VideoPinball-Atari2600 Viewpoint-Genesis Vigilante-Sms
VirtuaFighter-32x VolguardII-Nes WWFArcade-Genesis WaniWaniWorld-Genesis Wardner-Genesis Warpman-Nes WaynesWorld-Nes
WereBackADinosaursStory-Snes WhipRush-Genesis Widget-Nes WizardOfWor-Atari2600 WizardsAndWarriors-Nes WiznLiz-Genesis Wolfchild-Genesis
Wolfchild-Sms Wolfchild-Snes WolverineAdamantiumRage-Genesis WonderBoyInMonsterWorld-Sms WorldHeroes-Genesis WrathOfTheBlackManta-Nes WreckingCrew-Nes
XDRXDazedlyRay-Genesis XKaliber2097-Snes XMenMojoWorld-Sms Xenon2Megablast-Genesis Xenophobe-Nes XeviousTheAvenger-Nes Xexyz-Nes
YarsRevenge-Atari2600 YoukaiClub-Nes YoukaiDouchuuki-Nes YoungIndianaJonesChronicles-Nes Zanac-Nes Zaxxon-Atari2600 ZeroTheKamikazeSquirrel-Genesis
ZeroTheKamikazeSquirrel-Snes ZeroWing-Genesis ZombiesAteMyNeighbors-Snes ZoolNinjaOfTheNthDimension-Genesis ZoolNinjaOfTheNthDimension-Sms ZoolNinjaOfTheNthDimension-Snes
Figure 8: Diversity of our training dataset. Our curated dataset aggregates an extremely wide range of scenarios.
Appendix B Implementation Details
B.1 Architecture
The latent action autoencoder adopts a Transformer architecture with 500M parameters. It consists of 16 encoder blocks and 16 decoder blocks using 1024 channels and 16 attention heads. The dimension of the latent actions is 32. The autoregressive world model adopts a 3D UNet architecture following SVD (Blattmann et al., 2023), with 1.5B trainable parameters and a memory length of 6. The default input resolution for both models is $256\times 256$.
B.2 Training
The latent action autoencoder is trained for 200K steps from scratch with a batch size of 960. We employ the AdamW optimizer (Loshchilov & Hutter, 2019) with a learning rate of $2.5\times 10^{-5}$ and a weight decay of 0.01. The hyperparameter $\beta$ is set to $2\times 10^{-4}$ to achieve a good balance between representation capacity and context disentangling ability.
The autoregressive world model is trained for 80K steps with a batch size of 64 and a learning rate of $5\times 10^{-5}$ on 16 NVIDIA A100 GPUs. We adopt a cosine learning rate scheduler with 10K warmup steps. To enhance prediction quality, exponential moving average is also applied during pretraining. The noise augmentation level is randomly selected from a range of 0.0 to 0.7, with an interval of 0.1.
For both latent action autoencoder training and world model pretraining, we randomly jitter the brightness of input frames to augment generalization ability. For all frames with varying aspect ratios, we apply center cropping to avoid padding irrelevant content. We check the meta information of all datasets and downsample the videos to approximately a frame rate of 10 Hz for training.
B.3 Sampling
By default, we use 5 sampling steps with a classifier-free guidance scale of 1.05 to generate new frames. We do not add noise to the historical frames but condition the world model on an augmentation level of 0.1 as we find this improves the results. A timestep shifting strategy (Kong et al., 2024) is also applied to enhance generation quality.
B.4 Visual Planning with Model Predictive Control
We summarize our model predictive control process for visual planning as below:
  1. 1.
    Given the current observation and the image of the final state, where the environment is about to restart with a positive reward, we calculate the current reward as the cosine similarity between them in RGB space. The planning objective is defined as the maximum reward obtained along the planned trajectory.
  2. 2.
    At each iteration, we sample a population of $N$ action sequences, each with a length of $L$, from a distribution. The initial distribution is set to a uniform distribution over four actions (LEFT, DOWN, UP, RIGHT).
  3. 3.
    For each sampled action sequence, the world model is used to predict the resulting trajectory, and the reward is calculated for each trajectory.
  4. 4.
    The top $K$ action sequences with the highest rewards are selected, and we update the distribution by increasing the sampling probabilities of these selected actions.
  5. 5.
    A new set of $N$ action sequences is sampled from the updated distribution, and the process repeats for $i$ Cross-Entropy Method iterations.
  6. 6.
    After $i$ optimization iterations, the first $T$ action in the action sequence with the highest probability is executed in the environment. The planning terminates either when the final state is achieved or when the search limit is exceeded.
In practice, we use $i=2$ Cross-Entropy Method iterations. For each iteration, $N=100$ action sequences with a length of $L=15$ are sampled, and the best $K=10$ samples are selected to update the action sampling distribution. After the optimization procedure is done, the first $T=5$ actions are executed in the environment. We set the search limit to 20 steps.
Appendix C Additional Results
C.1 Action Transfer
Qualitative comparison. We qualitatively compare AdaWorld trained for 50K steps with other baselines in Figure 9. The results show that our approach is able to represent nuanced actions and exhibit strong transferability across contexts.
Visualizations. We attach more action transfer results in a number of environments in Figures 11, 12, 13, 14 and 15. Each sample is generated by transferring a latent action sequence with a length of 20.
Failure case study. We present typical failure cases of action transfer in Figure 20. AdaWorld does not always perfectly understand physics and dynamics. It also struggles to produce high-quality content during long-term rollouts or dramatic view shifts.
C.2 World Model Adaptation
To demonstrate the advantages of our approach, in Figure 10, we visualize some simulation results on the held-in test set using the finetuned models in Sec. 3.2.1. While the model pretrained with action-agnostic videos struggles to faithfully execute the action inputs, our model has rapidly acquired precise action controllability using the same number of steps. This underscores that our approach could serve as a better initialization method for world model adaptation compared to conventional methods.
C.3 Action Creation through Clustering
As mentioned in Sec. 2.3, AdaWorld can also easily create a flexible number of control options through latent action clustering. Specifically, we process our Procgen and Gym Retro training set using the latent action encoder to obtain the corresponding latent actions. To generate different control options, we apply K-means clustering to all latent actions, setting the number of clustering centers to the desired number of control options. To examine the controllability of varying actions derived with AdaWorld, we adopt the $\Delta$PSNR metric following Genie (Bruce et al., 2024). Table 7 shows the $\Delta$PSNR of the latent action decoder predictions. The larger the $\Delta$PSNR, the more the predictions are affected by the action conditions and therefore the world model is more controllable. The results in Table 7 demonstrate that the control options derived with AdaWorld represent distinct meanings and exhibit comparable controllability to the discrete counterpart, while the latter does not support a customizable number of actions, as it is fixed once trained.
Table 7: Action controllability evaluation. AdaWorld supports customizing different numbers of actions with strong controllability.
Method $\Delta$PSNR
4 5 6 7 8 9 10 11 12
Discrete cond. N/A N/A N/A N/A 6.47 N/A N/A N/A N/A
AdaWorld 5.67 5.15 7.28 8.23 6.26 7.32 6.07 6.68 6.53
Appendix D Something-Something v2 Categories for Action Transfer
In Sec. 3.1, we utilize the top-10 most frequently appearing categories from SSv2 for action transfer evaluation, including “Putting [something] on a surface”, “Moving [something] up”, “Pushing [something] from left to right”, “Moving [something] down”, “Covering [something] with [something]”, “Pushing [something] from right to left”, “Uncovering [something]”, “Taking [one of many similar things on the table]”, “Throwing [something]”, and “Putting [something] into [something]”.
Appendix E Selected Scenes for Visual Planning
To guarantee an effective evaluation, we conduct an exhaustive search using the ground truth environments to identify viable Procgen test scenes that can be completed in an acceptable number of steps by goal image matching. This results in a total of 120 scenes for our visual planning experiments in Sec. 3.2.2. In Figures 19, 16, 17 and 18, we illustrate the initial states of all selected scenes. The scenes from survival-oriented environments (e.g., BigFish and BossFight) are not considered in our evaluation.
Figure 9: Qualitative comparison of action transferability. AdaWorld can accurately identify the demonstrated action and precisely transfer it to another context, while the other baselines fall short in doing so.
Figure 10: Minecraft adaptation results after finetuning 800 steps with 100 samples for each action. Our approach efficiently achieves precise action controllability with minimal action-labeled data and finetuning, while the action-agnostic pretraining baseline fails to perform the correct actions.
Figure 11: Additional action transfer results by AdaWorld. Best viewed with zoom-in.
Figure 12: Additional action transfer results by AdaWorld. Best viewed with zoom-in.
Figure 13: Additional action transfer results by AdaWorld. Best viewed with zoom-in.
Figure 14: Additional action transfer results by AdaWorld. Best viewed with zoom-in.
Figure 15: Additional action transfer results by AdaWorld. Best viewed with zoom-in.
Figure 16: Heist scenes for planning evaluation. The figure illustrates the initial states of all selected scenes.
Figure 17: Jumper scenes for planning evaluation. The figure illustrates the initial states of all selected scenes.
Figure 18: Maze scenes for planning evaluation. The figure illustrates the initial states of all selected scenes.
Figure 19: CaveFlyer scenes for planning evaluation. The figure illustrates the initial states of all selected scenes.
Figure 20: Examples of failure cases. AdaWorld is by no means perfect in simulating real-world physics, dynamic agents, long-term rollouts, and significant view changes.